A 2000 kg car is accelerated uniformly from rest to a speed of 15 m/s in 10 s. What is the magnitude of the unbalanced force acting on the car?


1
Expert's answer
2020-12-17T09:13:39-0500

Let's first find the acceleration of the car from the kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{15\ \dfrac{m}{s}-0\ \dfrac{m}{s}}{10\ s}=1.5\ \dfrac{m}{s^2}."

Finally, we can find the magnitude of the unbalanced force acting on the car from the Newton's Second Law of Motion:


"F=ma=2000\ kg\cdot 1.5\ \dfrac{m}{s^2}=3000\ N."

Answer:

"F=3000\ N."